3/5 (4n+1)=23
I am having difficulty solving this problem. Due to the only common denominator I can find for 23 is 69.
I flipped the fraction and multiplied and I get:
20/3N + 5/3 = 115/3
- 5/3 -5/3
20/3N = 110/3
_____ _____
20/3 20/3
N=11/2 which is not correct
Then I tried multiplying everything by 69 which is the LCD if you flip everything and this did not work either. Can you help me solve this equation?
69(20/3N)+69(5/3)=69(23)
460N + 115=1587
- 115= -115
460N 1472
---- ------
460 460
N=16/5
This is nor correct either.

Answer by richwmiller(17219)

3/5*(4n+1)=23
(4n+1)=23*5/3
4n+1=115/3
Your big mistake was multiplying the left side by 5/3 twice.
5/3*3/5=1
4n=118/3
2n=59/3
n=28/3